OpenStudy (anonymous):

PreCalc: Evaluate the expression v ⋅ w Given the vectors: r = <5, -5, -2>; v = <2, -8, -8>; w = <-2, 6, -5>

OpenStudy (anonymous):

multiply the corresponding components then add them up

OpenStudy (anonymous):

@sourwing 2*-2=-4 -8*6=-48 -8*-5=40 -4 + -48 + 40= -12 this right?

OpenStudy (anonymous):

that is right